Three consecutive even integers are such that the square of the third is 76 more than the square of the second. Find the three integers.
Answer by checkley77(12844)

LET X, X+2 & X+4 BE THE THREE EVEN INTEGERS.
(X+4)^2=76+(X+2)^2
X^2+8X+16=76=X^2+4X+4
8X-4X=4+76-16
4X=64
X=64/4
X=16 ANS FOR THE SMALLEST INTEGER.
16=2=18 FOR THE MIDDLE INTYEGER.
16+4=20 FOR THE LARGEST INTEGER.
PROOF:
20^2=76+18^2
400=76+324
400=400